# Towards scalable robotic intervention of children with Autism Spectrum Disorder using LLMs

### v1

@@ VI-B Evaluation metrics
distilbert-base-uncased |
0.78 | 0.79 | 0.79 | 0.75 | 0.75 | 0.75 | 0.84 | 0.85 | 0.85 | |
BART |
roberta-base |
0.88 | 0.88 | 0.88 | 0.76 | 0.77 | 0.77 | 0.86 | 0.86 | 0.86 |
mnlimicrosoft/deberta-xlarge-mnli |
0.62 | 0.61 | 0.61 | 0.57 | 0.54 | 0.55 | 0.68 | 0.66 | 0.67 | |
distilbert-base-uncased |
0.71 | 0.76 | 0.73 | 0.70 | 0.74 | 0.72 | 0.80 | 0.83 | 0.81 | |
GPT-2 |
roberta-base |
0.85 | 0.87 | 0.86 | 0.84 | 0.85 | 0.84 | 0.90 | 0.91 | 0.90 |
mnlimicrosoft/deberta-xlarge-mnli |
0.53 | 0.57 | 0.55 | 0.49 | 0.53 | 0.51 | 0.69 | 0.71 | 0.70 |
VI-C Ablation study
This section compares using a BART pipeline for question and options generation vs an analogous GPT -2-based generation task. We evaluate the BERTscore for generated questions, options, and both for each test sample. As seen from Table I, our BART pipeline performs better on average for all of the generation tasks, at least when we use the Microsoft/deberta-xlarge-mnli model for calculating the BERTscore. This result contributes to our choice of using the BART pipeline because we use a Masked Language Modeling approach for it. Using masks makes this pipeline a supervised learning approach, which gives control over text generation.
Context |
Reference |
BART |
GPT-2 |
After several days of arguing with the boss, Bailey finally was exhausted. Bailey let it be. | Question: | Question: | Question: |
How would Bailey feel afterwards? |
What will Bailey want to do next? |
What will Bailey want to do next? |
Options: | Options: | Options: | |
A: ignored and defeated |
A: go home and rest |
A: sleep |
B: amenable |
B: get a new job |
B: break a sweat |
C: happy that the boss reconsidered |
C: go back to work |
C: quit the battle |
Cohen’s kappa (weighted) for BART pipeline |
R1 and R2 (0.705) |
R1 and R2 (0.639) |
R1 and R4 (0.693) |
R1 and R4 (0.787) |
R1 and R3 (0.72) |
R2 and R6 (0.680) |
R1 and R5 (0.85) |
R1 and R4 (0.70) |
R1 and R6 (0.80) |
R1 and R6 (0.634) |
R2 and R6 (0.85) |
R2 and R3 (0.715) |
R4 and R5 (0.87) |
R3 and R6 (0.8344) |
Cohen’s kappa (weighted) for GPT-2 pipeline | |||
R2 and R6 (0.64) |
An example comparing the BART and the GPT-2 pipeline for question and options generation can be seen in Table II. For a context taken from the SOCIALIQA dataset, we generate the text, evaluate it with the reference question and options from the dataset and calculate the score. We take five random examples from the SOCIALIQA dataset and ask six raters to score them between 0-10. Cohen’s kappa (weighted) was calculated to check for inter-rater reliability for each generation task (separately for question generation, options generation, and question + options generation) for each of the BART and GPT-3 pipelines. Table III shows the kappa values between the raters. We have only considered those values between the raters with a kappa higher than 0.6 (moderate agreement). Based on this threshold, the kappa values for the GPT-2 pipeline were only higher than 0.6 for rating . This low value can be attributed to fewer examples (five) considered for ratings. For comparing the BART and GPT pipelines based on the human evaluations for the whole text generation pipeline (), Figure 5 compares the values of the BERTscore with human evaluations for BART and GPT-2.
As evident from Figure 5, the BART pipeline performs better on the BERTscore metric, which is also consistent with human-evaluations. The human evaluation scores are very low for GPT-2 compared to the BERTscores since the human evaluations were performed on a very small sample of data with six raters.
VI-D Appropriateness for children with ASD
We finally concatenated our fine-tuned GPT-2 model from Section VI-A with the BART pipeline to generate the context, question, and options. An example from the generated texts for our GPT-2 + BART pipeline has been shown in Table IV. These texts were rated by human evaluators on two parameters:
Context |
Question |
Options |
Addison made Sydney’s bed for them. |
What will Sydney want to do next? |
A. thank Addison for helping |
B. get a new bed |
C. go to the store |
Q1: Is the social situation appropriate for children with ASD between the age range of 8-12 years (0-10 scale)?
Q2: Is there at least one appropriate answer from the options generated by the robot (0 or 1 scale)?
The scores for the first question were divided into binary scale (rater scores 0-7: 0 and 8-10: 1) for the inter-rater agreement between raters and 0-10 scale for finding the average of the scores. Four raters (R1, R2, R3, and R6) from Table III have clinical or teaching experience with children with ASD. The highest Cohen’s kappa was found between R2 and R6 (0.68). So, we calculated the average scores of these raters and the average score for Q1 was found to be 6.8 on a scale of 0-10. In addition, we found the percentage agreement between the raters for Q2 to be 0.95.
@@ VII Conclusion
In this paper, we attempt a scalable solution for a robotic intervention of children with ASD. We use an LLM pipeline with GPT-2 and BART to generate texts that the robot vocalises and can be used in a clinical setting in the presence of experts. The robot in this paper has been given the roles of a stimulator, prompter, and reinforcer, out of which the stimulator roles have autonomy in the text generation the robot uses. The automatic text generation is evaluated not only through the BERTscore metric but also through human evaluators. In addition, the appropriateness of the generated social situation using our LLM pipeline is also scored by experts with backgrounds in clinical intervention or teaching children with ASD.
As mentioned in the earlier sections, this work is the first of its kind to develop such a robot to automatically generate content during an intervention scenario for children with ASD. So, the scope of future work is immense. This system can be made more versatile in scenarios to be taught. For example, the LLM pipeline can be fine-tuned to generate content specific to a school setting, home setting, or even for making natural conversations, depending on the dataset used for fine-tuning. Therefore, the direction of our future work too lies in expanding the number of social situations for children with ASD and adapting to their specific needs during the intervention.

### v3

@@ V Results and Discussions
distilbert-base-uncased | 0.78 | 0.79 | 0.79 | 0.75 | 0.75 | 0.75 | 0.84 | 0.85 | 0.85 | |
BART | roberta-base | 0.88 | 0.88 | 0.88 | 0.76 | 0.77 | 0.77 | 0.86 | 0.86 | 0.86 |
mnlimicrosoft/deberta-xlarge-mnli | 0.62 | 0.61 | 0.61 | 0.57 | 0.54 | 0.55 | 0.68 | 0.66 | 0.67 | |
distilbert-base-uncased | 0.71 | 0.76 | 0.73 | 0.70 | 0.74 | 0.72 | 0.80 | 0.83 | 0.81 | |
GPT-2 | roberta-base | 0.85 | 0.87 | 0.86 | 0.84 | 0.85 | 0.84 | 0.90 | 0.91 | 0.90 |
mnlimicrosoft/deberta-xlarge-mnli | 0.53 | 0.57 | 0.55 | 0.49 | 0.53 | 0.51 | 0.69 | 0.71 | 0.70 |
As discussed previously, we have used the SOCIALIQA dataset for all of our generation tasks. We used the premium GPU from Google Colab with A100 GPUs for training. The dataset was split into train (75%), eval (15%), and test data (10%). Adam was chosen to be the optimizer. The SOCIALIQA dataset has one context, question, and three options for every data point. So, we fine-tuned GPT-2 for generating these contexts and then fed the output of the GPT-2 into our BART pipeline explained in the equation 2 to generate the questions and the options. We compared this pipeline with a similar pipeline, replacing BART with GPT-2 (both pipelines are shown in Figure 5).
V-A Context generation using GPT-2
For generating the context, we fine-tuned a GPT-2 model based on the contexts in the SOCIALIQA dataset. To demonstrate how similar the generated contexts were to the test set, we took one thousand GPT-2 generated texts (after fine-tuning) and the same number of data points from the test set. Further, we used the T-SNE plots to visualize how the test subset and the GPT-2 generated set overlap. We vectorized the generated texts and the subset of the test dataset using scikit-learn’s TfidfVectorizer to convert both the test and the generated datasets into a tf-idf feature matrix [55]. Once vectorized, the T-SNE plot was used to visualize the high-dimensional text data into two components. Since we wanted to compare how similar the GPT-2 generated texts and the actual test set, we just used the number of components to differentiate between them. As seen from Figure 7(a), the generated and the test set have a very large overlap. To characterize this overlap, we considered the data points of the scatter plot and used a Support Vector Classifier (SVC) with a radial basis kernel to see if the classifier could separate the data points into two distinct classes. Figure 7(b) and 7(c) show that the classification accuracy is 0.51, and the F1-score is around 0.55, too. This low F1-score signifies that the classifier could not very distinctly classify the values of the T-SNE plot (essentially the low-dimensional embeddings of the high-dimensional text data) [55]. The low scores on the classifying metrics signify that the text generated by fine-tuning GPT-2 generates data similar to the test set.
V-B Evaluation metrics
We use BERTscore to evaluate our question and option generation generation task [56]. We do this across generation of questions, options and, both of them combined. There are other metrics used in the literature like the BLEU score [57]. The reason for using BERTscore is because it uses cosine similarity instead of just making an n-gram based comparison between the actual text and the generated ones. In addition, for calculating the cosine similarities, it uses BERT embeddings between each generated token and every reference token [56]. This ensures not just exact phrase matching between the reference and the candidate text but also the contextual relevance between the two. Since there are different BERT based models to calculate the BERTscore from, we choose three of those models: 1) ‘distilbert-base-uncased’, 2) ‘roberta-base’ and, 3) ‘microsoft/deberta-large-mnlimicrosoft/deberta-xlarge-mnli’. Among these, microsoft/deberta-xlarge-mnli is currently the closest to human evaluations according to [58]. We evaluate the BERTscore for generated questions and options for each test sample. As seen from Table I, our BART pipeline performs better on average for all generation tasks, at least when we use the Microsoft/deberta-xlarge-mnli model for calculating the BERTscore.
V-C Self-reports by domain experts
In this section, we conduct statistical analysis on the self-reports collected from the domain experts about the robot session. As discussed previously, their responses were recorded on three surveys: NASA Task Load Index, Godspeed, and a self-curated survey for evaluating the appropriateness of our robotic session for children with ASD.
V-C1 NASA TLX
We used the NASA TLX to evaluate the domain experts’ mental workload (MWL) on six dimensions: mental demand (MD), physical demand (PD), temporal demand (TD), effort (Eff.), performance (Perf.), and frustration (Frus.) levels [59]. The experts were asked to fill the NASA TLX under two different circumstances. In one of the forms, they were asked to fill in the scores on a scale of 1-7 for the robot session. They were asked to do the same on a separate form for a classical perspective-taking session (without the robot). This was done to compare the MWL for the robot vs no-robot sessions.
Test for Normality: We tested the scores of the four domain experts for normality using Ryan-Joiner (similar to Shapiro-Wilk test) test for normality in Minitab. The difference in the ratings of , where for the robot vs. no robot sessions were used to check for normality as given in equation V-C1:
where is the domain expert’s rating on the NASA TLX for the robot session whereas represents the same for the no-robot session, and . The hypothesis for the Ryan-Joiner test is:
normal distribution | |||
normal distribution | (6) |
Since the p-value , for each of these dimensions (MD - Perf.), we accept the Null Hypothesis () that the distribution is normal.
Paired t-test: Since is normally distributed, it fulfills the assumption of the paired t-test. Hence, we compare the responses of the experts between the robot and no-robot sessions based on the following hypothesis:
where denotes the mean of the robot session and denotes the mean of the no-robot session.
As seen from Figure 9, there is no significant increase in the MWL of the domain experts when they are operating the robot. This is because the value for MS, PD, TD, Eff., and Frus., rejecting in equation V-C1. In addition to this observation, we can also see from Figure 9 that the value for Perf. and hence we reject from equation V-C1. However, as can be seen from Figure 9, the performance in the case of the robot session is higher than the no-robot session. The mean performance value was recorded as , compared to the in the no-robot session. Based on this comparison of means, we can conclude that the robot session does not demand any statistically significant additional MWL while being a scalable solution due to using LLMs. Moreover, the performance dimension of the NASA TLX shows a statistically significant increment in performance during the robot session compared to the no-robot session, which makes a strong case for domain experts to adopt such a system.
V-C2 Godspeed
To evaluate the perception of the domain experts about the robot, we asked them to fill in the Godspeed survey [60]. These perceptions are evaluated across five perceived features:
Perceived safety (PS)
anxious to relaxed
agitated to calm
quiescent to surprised
Anthropomorphism (AP)
fake to natural
machinelike to humanlike
unconscious to conscious
artificial to lifelike
moving rigidly to moving elegantly
Animacy (AN)
dead to alive
stagnant to lively
mechanical to organic
artificial to lifelike
inert to interactive
apathetic to responsive
Likeability (LK)
dislike to like
unfriendly to friendly
unkind to kind
unpleasant to pleasant
awful to nice
Perceived Intelligence (PI)
incompetent to competant
ignorant to knowledgeable
irresponsible to responsible
unintelligent to intelligent
foolish to sensible
Each of these perceived features, has different sub-categories. For example, perceived safety has three sub-categories. Each of these subcategories is rated by the domain experts on a scale of 1-5.
Test for normality: We follow the same hypothesis as in equation V-C1. The scores of the domain experts is represented by , where , , and represents the set of all subcategories of perceived feature , and the ideal scores are represented by .
Using the same Ryan-Joiner test for normality, and the hypothesis as shown in equation V-C1, we find value for , hence rejecting (distribution is normal).
Paired t-test:
where, is the mean value of of the perceived feature , and is the mean value of the ideal case i.e. . Figure 10 shows the results of the paired t-test. On comparing the mean values of the domain experts’ scores with the ideal case scenario, we reject for all subcategories for anthropomorphism and animacy and two subcategories for perceived intelligence. However, there was no significant statistical difference (value ) in the mean scores for every subcategory of perceived safety and likeability with the ideal case scenario. This means the robot was ideally perceived to be safe and likeable by the domain experts on the Godspeed surveys.
V-D Appropriateness for children with ASD
This survey was made specifically to take the domain experts’ opinions to evaluate the appropriateness of our human mediated robotic intervention. We ask them eight questions to include the aspects of our robotic session not covered in the previous two surveys:
Q1: How would you rate the roles of the robot as a initiator (content generator)?
Q2: How would you rate the roles of the robot as a prompter?
Q3: How would you rate the role of the robot as a reinforcer?
Q4: How appropriate was this entire session with the robot acting as a provoker, prompter, and reinforcer for children with ASD?
Q5: Would you prefer this human-mediated autonomous system as compared to a completely teleoperated system?
Q6: Did you find the robot reliable (with no technical glitches) through the entire duration of the session?
Q7: Would you prefer this human-mediated autonomous system in a therapeutic or educational settings?
Q8: Did you find the Graphical Use Interface (GUI) helpful to operate the robot?
For each question, Qi from , we ask the domain experts to give their ratings on a scale of 1-5.
Test for normality: For a question , the scores of the domain experts were compared against the ideal scores .
The hypothesis used in equation V-C1 were used to conduct the Ryan-Joiner test for normality. It was found that for all , the value , hence accepting H0 that the distribution is normal.
Paired t-test:
where is the mean of the domain experts’ responses on and is the mean for the ideal case scenario . From Figure 12, it can be seen that the roles of the robot for the entire session acting as an initiator, prompter, and reinforcer (Q4) were statistically no different from the ideal case scenario (value ). In addition, the roles of the robot as a prompter and reinforcer individually too were not statistically significant from the ideal case scenario (value ). Further, based on the same rational of comparing the means of the domain experts scores to the ideal case scenario, the domain experts preferred our human-mediated autonomous system compared to a completely teleoperated system (Q5). This is because our human mediated system was reliable during the time the robot was being operated by the domain expert. It faced practically no technical glitches that would hinder the progress of the session (Q6). This can be seen from (value ) for Q6. To make the system more user friendly, the GUI as shown in Figure 4 was also considered very helpful to operate the robot (Q8 and Figure 12).
The scores for Q1 and Q7 came out to be statistically different from the ideal case scenario. However, for Q1, an important point to note is that the power of the paired t-test is 0.6433 (see Figure 13). So, because of this low power, it is difficult to reliably reject the null hypothesis in equation V-D. But for Q7, the power comes out to be 0.917. Although for Q7, the mean is statistically different from the ideal case scenario, the mean for Q7 is 4.00 which is more than half the length of the 5-point likert scale. In addition, the low score of Q7 can be attributed to the use of a dataset (SOCIALIQA) (used fine-tuning our LLM approach) that was not built keeping in mind children with ASD. However, this human mediated autonomous system mitigates this effect since the authority to override the robot lies with the domain expert (explained in Section III-A).
VI Limitations and Future work
Although we could make conclusions based on the domain experts’ responses on different surveys, it will be more helpful to collect data from more such experts. This would probably help us get different results, such as in Figure 12. In Figure 12, except for Q1 and Q7, all other responses were statistically similar to the ideal cases. Furthermore, having more number of domain experts to rate might impact the results of the paired t-test shown in Figure 10.
Beyond the self reports, we would like to record naturalistic conversations between domain experts and children with ASD. This will help us achieve multiple goals. First, we can extend our current work by fine-tuning LLMs into making natural conversations with the child with ASD. Second, we can create a dataset of the communications recorded between the domain experts and the robot which would help us tailor our LLM models to specific scenarios, thus personalizing this robot experience for children with ASD.
@@ VII Conclusion
In this paper, we develop an intervention scheme for children with ASD and introduce autonomy in one of the elements of the pipeline, i.e. content generation. This helps make the approach more scalable in the sense that an LLM pipeline can be fine-tuned to generate contents specific to a situation, or to a therapeutic goal or outcome in mind. The NAO robot used in this paper acts as a initiator, stimulator, and a reinforcer during the intervention. This LLM pipeline contributes to the autonomy in the content generation for the initiator and the prompter roles of the robot. The decision on which role the robot should take up during the intervention lies with the human, hence contributing the human-mediated autonomy of the system. This gives control to the domain expert on what contents generated by the LLM pipeline is appropriate for the robot to speak during an intervention of children with ASD.
In addition to developing a technical pipline for a human-mediated autonomous system, we also conducted a user study to collect the opinion of domain experts. We do that by conducting a ten minute session with the robot, domain expert, and an actor participant. We collect the opinions on the domain experts based on three surveys: NASA TLX, Godspeed, and an appropriateness survey. We perform statistical tests to compare the responses of the domain experts and report it in Section V-C.
VIII Ethics approval
This study was approved by the University’s Institute Review Board (IRB). All the participants gave informed consent, and they had an option to opt out of the study at any stage.